OpenStudy (anonymous):

√ 108 - 3 √ 50 + 5 √ 12 + 4 √ 18

OpenStudy (anonymous):

first of all...simplify \[\sqrt{108}\], \[\sqrt{50}\], \[\sqrt{12} and \sqrt{18}\]

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

16 √ 3 - 3 √ 2

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\[\sqrt{108} is 6\sqrt{3}....\]

OpenStudy (anonymous):

\[-3\sqrt{50} is -3*5\sqrt{2}...which is 15\sqrt{2}\]

OpenStudy (anonymous):

\[5\sqrt{12} is 5*2\sqrt{3}...so \it's 10\sqrt{3}\]

OpenStudy (anonymous):

\[4\sqrt{18} is 4*3\sqrt{2}...which is 12\sqrt{2}\]

OpenStudy (anonymous):

\[now find the answer for 6\sqrt{3}-15\sqrt{2}-10\sqrt{3}+12\sqrt{2}...\]